OpenStudy (anonymous):

5x – 7(x + 1) > –9 + 3

OpenStudy (anonymous):

\[5x-7(x+1)=5x-7x-7=-2x-7\] is a start

OpenStudy (anonymous):

and \(-9+3=-6\) so start with \[-2x-7>-6\]

OpenStudy (anonymous):

add \(7\) and get \(-2x>1\) divide by \(-2\) and change the sign this time to finish with \[x<-\frac{1}{2}\]